an investor deposits $5250 in a certificate of deposit which pays 7.35% and compounds weekly. What will be the balance in seven (7) years.

Please show formula! Thanks.
Answer #1

Formula is C =P *(1+r/n)^tn

where C is the total returns, P is the initial deposit amount, r is the annual interest rate in decimal form, n is number of compounding periods per year (for example, 12 for monthly compounding) and t is the numbers of years the money is invested for.

C=5250*(1+(7.35%/52)^52*7

C=5250*(1+0.001413)^52*7

C=5250*(1.0014)^364

C=5250*1.672194

C=8779.02